OpenStudy (anonymous):

forgottn how to do this: 42d+1=20d-1

OpenStudy (anonymous):

collect your like tersm on each side and then solve for the variable. What do you get when you collect like terms?

OpenStudy (anonymous):

22d = 0

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Not zero, check your right side. You need to bring the +1 over from the left side, crossing the bridge over the equal sign, forcing it to change signs. You are left with -1-1 which is not zero but -2. No wsolve for the variable.

OpenStudy (anonymous):

okay so i done that ... but i dont think -2 can be divided by 22..so is that the answer

OpenStudy (anonymous):

You will now have: \[42d+1=20d\] \[42d-20d=-1-1\] \[22d=-2\] \[d=\frac{ -2 }{ 22 }\] d=-1/11

OpenStudy (anonymous):

If you want you can turn the fraction into a decimal but the fraction is acceptable to leave in its lowest terms. Yes, you can distribute the negative sign along with the value as you divide right through.

OpenStudy (anonymous):

oh okay thanks man

OpenStudy (anonymous):

np